Two trains are moving in opposite directions at the speed of 50km/hr and 70km/hr. Their lengths are 150m and 100m. The time required for their crossing each other is?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Two trains are moving in opposite directions at the speed of 50km/hr and 70km/hr. Their lengths are 150m and 100m. The time required for their crossing each other is?
------------
Their speeds are added when going in opposite directions = 120 km/hr
"crossing" is the sum of the lengths = 250 m = 0.25 km
-----------
d = r*t
t = d/r = 0.25/120 hours

Answer by stanbon(75887)

Two trains are moving in opposite directions at the speed of 50km/hr and 70km/hr. Their lengths are 150m and 100m. The time required for their crossing each other is?
----
Their crossing rate is 50+70 = 120km/hr
------
Their combined lengths is 150 + 100 m = 250 m
------
time = distance/rate = 250 m/120000m/hr = 0.0021 hr = 0.125 minutes = 7.5 sec
-----------
Cheers,
Stan H.
----------